Find the length of a rectangular lot with a perimeter of 108 meters if teh length is 8 meters more than the width.
Answer by stanbon(75887)

Find the length of a rectangular lot with a perimeter of 108 meters if teh length is 8 meters more than the width.
----------------------------------
Perimeter = 2(W + L)
108 = 2(W + W+8)
54 = 2W+8
2W = 46
W = 23 meters (width)
w+8 = 31 meters (length)
============================
Cheers,
Stan H.